# In-Context Learning for Seismic Data Processing

### v1

@@ 1 Introduction
Figure 1: Comparison of supervised learning, visual prompting, and in-context learning. Supervised Learning predicts $Y^{*}$ based solely on $X$. In-context learningpredicts $Y^{*}$ based on $X$ and a support set $S$ containing CDPsfrom the same seismic line as $X$ and labels obtained via an arbitrary demultiple process (e.g. Radon).
@@ 2.3 Synthetic Dataset and Training
   $X\leftarrow X^{\text{Train}}[n,i_{0}]$ $\triangleright$ First index $i_{0}$ is for the input $X$ and output $Y$.
   $Y\leftarrow Y^{\text{Train}}[n,i_{0}]$
@@
    $V[s-1,0]\leftarrow X^{\text{Train}}[n,i_{s}]$
    $V[s-1,1]\leftarrow Y^{\text{Train}}[n,i_{s}]$
@@
   $X,Y,V\leftarrow\text{RandomWhiteNoise}(X,Y,V)$ $\triangleright$ Add random white noise to all the gathers as well as labels.
   $X,Y,V\leftarrow\text{NormalizePerImage}(X,Y,V)$ $\triangleright$ Normalize each gather and label by the mean and standard deviation of the gather.
   $X,Y,V\leftarrow\text{RandomlyReplaceLabelWithInput}(X,Y,V)$ $\triangleright$ For a specified percentage of steps replace the label with the input, as well as the prompt-labels with the prompts.
@@
   $Y^{*}\leftarrow f^{\text{ICL}}_{\theta}\left(X,V\right)$ $\triangleright$ Prediction.
   $\ell\leftarrow\mathcal{L}_{1}(Y^{*},Y)$ $\triangleright$ Compute loss.

### v2

@@ 1 Introduction
Figure 1: Comparison of supervised learning and in-context learning. Supervised Learning predicts $\mathbf{Y}^{*}$ based solely on $\mathbf{X}$. In-context learningpredicts $\mathbf{Y}^{*}$ based on $\mathbf{X}$ and a support set $\mathcal{V}$ containing CDPsfrom the same seismic line as $\mathbf{X}$ and labels obtained via an arbitrary demultiple process (e.g. Radon).
@@ 2.3 Synthetic Dataset and Training
   $\mathbf{X}\leftarrow\mathbf{X}^{\text{Train}}[n,i_{0}]$ $\triangleright$ First index $i_{0}$ is for the input $\mathbf{X}$ and output $\mathbf{Y}$.
   $\mathbf{Y}\leftarrow\mathbf{Y}^{\text{Train}}[n,i_{0}]$
@@
    $\mathbf{V}[s-1,0]\leftarrow\mathbf{X}^{\text{Train}}[n,i_{s}]$
    $\mathbf{V}[s-1,1]\leftarrow\mathbf{Y}^{\text{Train}}[n,i_{s}]$
@@
   $\mathbf{X},\mathbf{Y},\mathbf{V}\leftarrow\text{RandomWhiteNoise}(\mathbf{X},\mathbf{Y},\mathbf{V})$ $\triangleright$ Add random white noise to all the gathers as well as labels.
   $\mathbf{X},\mathbf{Y},\mathbf{V}\leftarrow\text{NormalizePerImage}(\mathbf{X},\mathbf{Y},\mathbf{V})$ $\triangleright$ Normalize each gather and label by the mean and standard deviation of the gather.
   $\mathbf{X},\mathbf{Y},\mathbf{V}\leftarrow\text{RandomlyReplaceLabel}(\mathbf{X},\mathbf{Y},\mathbf{V})$ $\triangleright$ For a specified percentage of steps replace the label with the input, as well as the prompt-labels with the prompts.
@@
   $\mathbf{Y}^{*}\leftarrow f^{\text{ICL}}_{\theta}\left(\mathbf{X},\mathbf{V}\right)$ $\triangleright$ Prediction.
   $\ell\leftarrow\mathcal{L}_{1}(\mathbf{Y}^{*},\mathbf{Y})$ $\triangleright$ Compute loss.